# Carl Størmer and his Numbers

### v1

@@ 1. Introduction and Motivation
For $m\in\mathbb{N}$ and positive integers $q_{1},q_{2},\cdots,q_{m}\in\mathbb{N}$, the continuant $q_{1},q_{2},\cdots,q_{m}]$ of length $m$ is defined by $q_{1}]=q_{1}$ and, for $m>1$, it is the determinant
@@
$s=[q_{1},q_{2},\cdots,q_{n}]\quad\text{and}\quad r=[q_{2},q_{3},\cdots,q_{n}].$
@@ 2. Størmer Numbers
where $r$ is a non-negative integer, each $r_{j}\in\mathbb{N},$ and each $p_{j}$ is an odd prime number with $p_{1}<p_{2}<\ldots<p_{m-1}<p_{m}.$
@@
Suppose $n=x_{0}^{2}+1$ has the prime factorization given in (2.5) where we assume the primes satisfy $p_{1}<p_{2}<\ldots<p_{m}$. Then there exists a prime $p$ for which $x_{0}$ is the Størmer number if and only if $2x_{0}+1\leq p_{m}$. Moreover, if this condition is met, then $p=p_{m}$ and $S(p_{m})=x_{0}.$ Otherwise, $x_{0}$ is not the Størmer number for any prime $p\equiv 1\bmod 4.$
@@ 4. A Biographical Sketch of Carl Størmer
Størmer entered the University of Christiana (Oslo’s previous name) in 1892 and obtained his candidates’ degree (similar to a Ph.D. degree) in 1897. By this time, Størmer had written several mathematical articles, primarily in number theory, as well as a number of short notes in botany. From 1898-1900, Størmer studied at the Sorbonne under mathematical giants Émile Picard, Henri Poincaré, Camille Jordan, Gaston Darboux, and Édouard Goursat. His output of mathematical papers continued with twelve papers on series, number theory, and the theory of functions between 1896 and 1902. He worked with Sylow and Holst to produce two volumes in 1902 to celebrate the centenary of Niels Henrik Abel’s (1802-1829) birth.
@@
In 1903, Størmer was appointed professor of pure mathematics at the University of Oslo, a position he held until 1946. In 1934, Størmer published a paper in the Norwegian Mathematical Society entitled ‘Ramanujan - A Remarkable Genius’. At the insistence of fellow Norwegian Atle Selberg, a leading authority in number theory, this paper was reprinted in the same journal in 1989. Selberg was a teenager in 1934 and remarked that Størmer’s paper on Ramanujan contributed to arousing his interest in mathematics.

### v2

@@ 1. Introduction and Motivation
For $m\in\mathbb{N}$ and positive integers $q_{1},q_{2},\ldots,q_{m}\in\mathbb{N}$, the continuant $q_{1},q_{2},\ldots,q_{m}]$ of length $m$ is defined by $q_{1}]=q_{1}$ and, for $m>1$, it is the determinant
@@
$s=[q_{1},q_{2},\ldots,q_{n}]\quad\text{and}\quad r=[q_{2},q_{3},\ldots,q_{n}].$
@@ 2. Størmer Numbers
where $r$ is a non-negative integer, each $r_{j}\in\mathbb{N},$ and each $p_{j}$ is an odd prime number with $p_{1}<p_{2}<\cdots<p_{m-1}<p_{m}.$
@@
Suppose $n=x_{0}^{2}+1$ has the prime factorization given in (2.5) where we assume the primes satisfy $p_{1}<p_{2}<\cdots<p_{m}$. Then there exists a prime $p$ for which $x_{0}$ is the Størmer number if and only if $2x_{0}+1\leq p_{m}$. Moreover, if this condition is met, then $p=p_{m}$ and $S(p_{m})=x_{0}.$ Otherwise, $x_{0}$ is not the Størmer number for any prime $p\equiv 1\bmod 4.$
@@ 4. A Biographical Sketch of Carl Størmer
Størmer entered the University of Christiania (Oslo’s previous name) in 1892 and obtained his candidates’ degree (similar to a Ph.D. degree) in 1897. By this time, Størmer had written several mathematical articles, primarily in number theory, as well as a number of short notes in botany. From 1898-1900, Størmer studied at the Sorbonne under mathematical giants Émile Picard, Henri Poincaré, Camille Jordan, Gaston Darboux, and Édouard Goursat. His output of mathematical papers continued with twelve papers on series, number theory, and the theory of functions between 1896 and 1902. He worked with Sylow and Holst to produce two volumes in 1902 to celebrate the centenary of Niels Henrik Abel’s (1802-1829) birth.
@@
In 1903, Størmer was appointed professor of pure mathematics at the University of Oslo, a position he held until 1946. In 1934, Størmer published a paper in the Norwegian Mathematical Society entitled ‘Ramanujan – A Remarkable Genius’. At the insistence of fellow Norwegian Atle Selberg, a leading authority in number theory, this paper was reprinted in the same journal in 1989. Selberg was a teenager in 1934 and remarked that Størmer’s paper on Ramanujan contributed to arousing his interest in mathematics.